OpenStudy (anonymous):

3- (x-1)=x I got 3= 2x do i just dived that

hartnn (hartnn):

how did u get 3=2x try again,its incorrect.

OpenStudy (anonymous):

Well do you multiply 3 and x? or do you just do 3-1x=x

OpenStudy (anonymous):

that it: 3+1-x=x 4=2x x=2

hartnn (hartnn):

1st add (x-1) to both sides,u get 3=x+x-1 now x+x=2x 3=2x-1 add 1 to both sides,what u get?

OpenStudy (anonymous):

2:)

hartnn (hartnn):

correct :) x=2